11. XYZ Company is a retailer of widgets. XYZ pays $4.75 for each widget and sells them for $8.00. Monthly fixed costs are $26,000. The widget cost is the only variable cost.

a. What is the contribution margin per unit?
b. What is the break-even point in units?
c. How many units will XYZ need to sell to earn target profit of $13,000?
Answer #1

1) The formula to calculate contribution margin per unit = unit price
Answer #2
Particulars ($)
1 Variable cost              4.75
2 Selling orice of widget                    8
3 Fixed Cost          26,000
4 Contribution Margin (2-1)              3.25
5 Break Even Point (in units) [3/(2-1)]            8,000
6 Units (Fixed expenses + Target profit) / Unit contribution margin       12,000
Answer #3

Cost= $4.75

Sale= $8.00

Fixed Cost= 26000

a. Contribution Margin Per Unit=

Contribution margin is defined as revenues minus variable expenses= 8-4.75= $3.25

b. Breakeven is where Revenue-Cost= 0, ie. Profit is Zero.

Let x be the number of Units

26000+4.75x= 8x

X= 26000/3.25= 8000 Units

c. 13000= 8x-26000-4.75x

39000= 3.25x

X= 39000/3.25= 12000 Units